OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80 A. an = -5 • (-4)n B. an = -5(-4)n-1 C. an = -4(-5)n-1 D. an = -5 • (4)n

OpenStudy (anonymous):

Just waiting for some help.........

OpenStudy (anonymous):

\(\huge {a _{n} = a _{1}.r^{n-1}}\)

OpenStudy (anonymous):

see, \(\huge a _{1} = -5 \) find \(\huge r\) :))

OpenStudy (anonymous):

\(\huge r=\frac{ a _{2} }{ a _{1} }\)

OpenStudy (anonymous):

So that would be \[\frac{ 5^{_{2}} }{ 5_{1} }\]

OpenStudy (anonymous):

no, see \(\huge a _{2} = 20 \) :D

OpenStudy (anonymous):

So \[\frac{ 20 }{ 10 } then?\]

OpenStudy (anonymous):

\(\huge a _{1} = -5 \) \(\huge r=\frac{ 20 }{ -5 }\)

OpenStudy (anonymous):

So we got r, where do I go from there?

OpenStudy (anonymous):

\(\huge {a _{n} = a _{1}.r^{n-1}}\) you will find \(\huge a _{n}\) by your \(\huge r\) :))

OpenStudy (anonymous):

I don't know how to solve that.

OpenStudy (anonymous):

whats your \(\huge r \) ?

OpenStudy (anonymous):

20/-5

OpenStudy (anonymous):

Whats the result of it?

OpenStudy (anonymous):

i mean 20/-5= ?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

Yes, then an= -5.(-4)^n-1

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

hope you understand :)

OpenStudy (texaschic101):

an = a1 * r ^ n - 1 r = the difference between each term....what you multiply by to get the next number a1 = the first number